OpenStudy (hba):

Integrate :

OpenStudy (hba):

\[\int\limits_{}^{} \frac{ x }{ \sqrt{4+x ^{2}} }dx=\]

hartnn (hartnn):

put x^2+4=u

OpenStudy (hba):

Ok so i substitute it, What next ??

OpenStudy (hba):

\[x=\sqrt{u-4}\]

hartnn (hartnn):

dx= ?

OpenStudy (hba):

ohh i should derivate it

hartnn (hartnn):

lol. no , don't try to find x....

hartnn (hartnn):

differentiate this : x^2+4=u

OpenStudy (hba):

dx=du/2x

hartnn (hartnn):

that is xdx = du/2 ok?

OpenStudy (hba):

Yeah right

hartnn (hartnn):

so u get int of 1/ root u du/2 right ?

OpenStudy (hba):

\[\int\limits_{}^{}\frac{ du }{ \sqrt{u} 2}\]

hartnn (hartnn):

yup, so 1/root u = u^{what} ?

OpenStudy (hba):

u^-1/2

hartnn (hartnn):

can u integrate that?

OpenStudy (hba):

2u^1/2

hartnn (hartnn):

isn't it just root u ?

OpenStudy (hba):

yeah If we cancel it Only root u is left what next ??

hartnn (hartnn):

resubstitute u = ....

hartnn (hartnn):

add +c in the end

OpenStudy (hba):

\[\sqrt{4+x ^{2}}+c\]

hartnn (hartnn):

that is correct.

OpenStudy (hba):

Thankyou very Much :)

hartnn (hartnn):

welcome ^_^